# Executable Knowledge Graphs for Replicating AI Research

### v1

@@ Abstract
Replicating AI research is a crucial yet challenging task for large language model (LLM) agents. Existing approaches often struggle to generate executable code, primarily due to insufficient background knowledge and the limitations of retrieval-augmented generation (RAG) methods, which fail to capture latent technical details hidden in referenced papers. Furthermore, previous approaches tend to overlook valuable implementation-level code signals and lack structured knowledge representations that support multi-granular retrieval and reuse. To overcome these challenges, we propose Executable Knowledge Graphs (xKG), a modular and pluggable knowledge base that automatically integrates technical insights, code snippets, and domain-specific knowledge extracted from scientific literature. When integrated into three agent frameworks with two different LLMs, xKG shows substantial performance gains (10.9% with o3-mini) on PaperBench, demonstrating its effectiveness as a general and extensible solution for automated AI research replication [^1] https://github.com/zjunlp/xKG..
Executable Knowledge Graphs for Replicating AI Research
@@ 1 Introduction
Why Executable Knowledge Graphs? Existing attempts Tang et al. (2025); Ou et al. (2025) to convert papers into knowledge bases show promising signs but often stop at shallow scaffolding rather than delivering rigorous, reproducible implementations. Three key issues limit agent-driven reproduction: (1) most approaches fail to extract deeper technical insights hidden in cited references and background literature; (2) they overlook practical signals embedded in concrete code implementations; and (3) the lack of a structured, unified representation prevents effective retrieval, composition, and reuse of scientific concepts and their executable components Hua et al. (2025).
To address these gaps, we propose the Executable Knowledge Graph (xKG), a novel knowledge representation that fuses textual paper knowledge with its corresponding executable code snippets. Unlike conventional KG, xKG captures both conceptual relations and runnable components, enabling agents to retrieve, reason about, and assemble the precise artifacts needed for faithful reproduction. We evaluate xKG by integrating it into three distinct agent frameworks—BasicAgent, IterativeAgent, and PaperCoder. Our experiments on PaperBench Starace et al. (2025) demonstrate consistent and significant performance gains. The design of xKG is modular and extensible, facilitating its adoption and expansion across diverse research domains.
@@ 2 Executable Knowledge Graphs
Figure 1: xKG is constructed automatically from arXiv papers and GitHub repositories (Examples at Appendix D).
2.1 Preliminary
We define the paper reproduction task as generating an executable code repository $R$ from a paper $P$, modeled as $R=\mathcal{A}(P)$, where $\mathcal{A}$ is an agent. The primary benchmark for this task evaluates the functional correctness of $R$ against an evaluation rubric $\mathcal{T}$. A final Replication Score, $S=\mathcal{E}(R,\mathcal{T})$, quantifies the weighted proportion of criteria met.
@@ 2.2 Design Formulation
We model xKG as a hierarchical, multi-relational graph $\mathrm{\textsc{xKG}}=(\mathcal{N},\mathcal{E})$, which is composed of various node types and edge types defined as:
@@
We first define the three types of nodes to capture knowledge at different granularities:
@@
  • Technique Node ($n_{t}$): A self-contained academic concept $n_{t}=(D_{t},\{n^{\prime}_{t}\}_{k})$ with its definition $D_{t}$ and sub-nodes $\{n^{\prime}_{t}\}_{k}$, ranging from a complete framework to a reusable component.
@@
These nodes are then linked by the following two primary types of edges:
@@
  • Implementation Edge ($e_{\text{impl}}$): A directed edge $(n_{t},n_{c})$ linking a technique node to its code implementation.
Through the above design, it becomes possible to link specific techniques (e.g., “Contrast-Consistent Search” in Burns et al. (2022), see Figure 5) from papers, as well as associate these techniques with code snippets, yielding more precise knowledge.
2.3 Executable Knowledge Graph Construction
2.3.1 Corpus Curation
Our corpus curation strategy is a fully automated, paper-centric pipeline designed for scalability. For each paper targeted for reproduction (papers in PaperBench Starace et al. (2025)), we employ o4-mini to identify its core techniques, which drive a two-pronged collection process. Note that we strictly do NOT use the GitHub repositories or third-party reproduction repositories listed in PaperBench’s blacklist to avoid any risk of data leakage. We first perform reference-based selection, expanding the corpus by filtering each paper’s references and retaining the top five most valuable works, ranked by their technical contribution and overlap. Next, we conduct technique-based retrieval, using the top-contributing techniques as keywords to retrieve additional papers from the web. All retrieved papers are processed to fetch their sources from arXiv and then identify the associated GitHub. A rule-based filter is applied to retain papers with official repositories, resulting in the curated corpus of paper–repository pairs.
@@ 2.3.1 Corpus Curation
 vanilla ++xKG  vanilla ++xKG  vanilla ++xKG  vanilla ++xKG  vanilla ++xKG  vanilla ++xKG
BasicAgent o3-mini 12.96 $37.22_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+24.26}}$ 22.63 $27.26_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+4.63}}$ 18.24 $20.82_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+2.58}}$ 20.82 $22.86_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+2.04}}$ 14.82 $14.67_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-0.15}}$ 17.89 $24.57_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+6.68}}$
DS-R1 33.05 $39.14_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+6.09}}$ 40.55 $39.14_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-1.41}}$ 17.22 $24.49_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+7.27}}$ 31.56 $33.97_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+2.41}}$ 17.08 $21.38_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+4.30}}$ 27.89 $31.62_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+3.73}}$
IterativeAgent o3-mini 22.22 $43.70_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+21.48}}$ 21.38 $36.28_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+14.90}}$ 28.77 $23.91_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-4.86}}$ 31.28 $29.15_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-2.13}}$ 19.35 $26.50_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+7.15}}$ 24.60 $31.91_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+7.31}}$
DS-R1 16.20 $47.40_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+31.20}}$ 31.19 $31.78_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+0.59}}$ 31.09 $26.57_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-4.52}}$ 35.30 $38.44_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+3.14}}$ 21.32 $31.89_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+10.57}}$ 27.02 $35.22_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+8.20}}$
PaperCoder o3-mini 23.15 $46.48_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+23.33}}$ 45.70 $53.99_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+8.29}}$ 52.48 $52.08_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-0.40}}$ 50.37 $63.13_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+12.76}}$ 39.84 $50.36_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+10.52}}$ 42.31 $53.21_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+10.90}}$
DS-R1 43.24 $49.26_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+6.02}}$ 43.26 $59.19_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+15.93}}$ 51.18 $73.03_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+21.85}}$ 61.12 $60.68_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-0.44}}$ 62.37 $59.53_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-2.84}}$ 52.23 $60.34_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+8.11}}$
@@
Table 1: Main results on PaperBench Code-Dev. We evaluate on the official lite subset of PaperBench, consisting of five papers: MU-DPO, TTA-FP, One-SBI, CFG, and FRE (details in Table 3). Results are reported using the Replication Score (%) metric with o3-mini as evaluator. All scores are shown as best@3 to mitigate task stochasticity and tool-related failures.
2.3.2 Hierarchical Graph Construction
@@
Based on the corpus obtained above, we then proceed to construct the xKG, including the following three automated steps:
@@
  • Step 1: Technique Extraction. We first use o4-mini to deconstruct the paper’s methodology into a preliminary hierarchical tree of Technique Nodes $\mathcal{N}_{T}$ linked by Structural Edges $e_{\text{struct}}$. Subsequently, we utilize RAG [^2] We employ text-embedding-3-small throughout all stages of xKG construction. (treating the paper as a document) to enrich each node by retrieving relevant text from the paper, which is then synthesized into a comprehensive definition $D_{t}$. This step yields a set of detailed yet unverified techniques that may contain noise.
@@
  • Step 2: Code Modularization. For each technique $n_{t}$, its definition is used as a query to retrieve relevant code snippets, following the similar RAG-based procedure (treating the code as a document) as in Step 1. We then employ o4-mini to synthesize these snippets into a candidate Code Node $n_{c}$, which includes the implementation $\sigma$, a test script $\tau$, and accompanying documentation $\delta$. This candidate node is then organized in a modular fashion and subjected to an iterative self-debugging loop to verify the executability of each module, ultimately producing a set of fully executable Code Nodes $\mathcal{N}_{c}$ along with their associated Implementation Edges $e_{\text{impl}}$.
@@
Finally, we construct xKG from 42 curated papers, totaling 591,145 tokens. We aim to automate this process to enable knowledge scaling.
2.4 Using Executable Knowledge Graphs
In a practical reproduction workflow, a LLM agent can use xKG at two critical stages. For high-level planning, the agent fetches the target paper’s Paper Node (without all Code Nodes) to grasp its core techniques and overall structure. During low-level implementation, the agent queries xKG for semantically relevant (Technique, Code) pairs to aid in specific functionalities. These two steps can be supplied either as callable tools for ReAct-style agents or as pluggable components of fixed-workflow agents. Crucially, to combat knowledge noise, all retrieved candidates are processed by a final LLM-based Verifier (o4-mini). This verifier acts as a critical quality gate, filtering, reranking, and refining the results to ensure that the retrieved knowledge is highly relevant and implementable.

### v2

# What Makes AI Research Replicable? Executable Knowledge Graphs as Scientific Knowledge Representations

@@ Abstract
Replicating AI research is a crucial yet challenging task for large language model (LLM) agents. Existing approaches often struggle to generate executable code, primarily due to insufficient background knowledge and the limitations of retrieval-augmented generation (RAG) methods, which fail to capture latent technical details hidden in referenced papers. Furthermore, previous approaches tend to overlook valuable implementation-level code signals and lack structured knowledge representations that support multi-granular retrieval and reuse. To overcome these challenges, we propose Executable Knowledge Graphs (xKG), a pluggable, paper-centric knowledge base that automatically integrates code snippets and technical insights extracted from scientific literature. When integrated into three agent frameworks with two different LLMs, xKG shows substantial performance gains (10.9% with o3-mini) on PaperBench, demonstrating its effectiveness as a general and extensible solution for automated AI research replication [^1] https://github.com/zjunlp/xKG..
What Makes AI Research Replicable?
Executable Knowledge Graphs as Scientific Knowledge Representations
@@ 1 Introduction
Why Executable Knowledge Graphs? AI research is hard to replicate and reuse because its knowledge is implicit and fragmented across text, code, and configuration. To address this challenge, we propose the Executable Knowledge Graph (xKG), a novel, paper-centric knowledge base designed to externalize latent scientific knowledge into a verifiable, executable representation.
Our xKG overcomes the limitations of existing attempts Tang et al. (2025); Ou et al. (2025), which often stop at coarse-grained knowledge reuse, by modeling scientific literature as hierarchical graphs that ground academic concepts in executable code. Unlike conventional KGs, xKG captures both conceptual relations and executable components, enabling agents to assemble the precise artifacts needed for faithful reproduction. We evaluate xKG by integrating it into three distinct agent frameworks: BasicAgent, IterativeAgent, and PaperCoder. Experiments on PaperBench Starace et al. (2025) demonstrate consistent and significant performance gains. Our xKG is built on a fully automated paper-aware pipeline, updatable to stay current with advancing research.
@@ 2 Executable Knowledge Graphs
Figure 1: The Paper-Centric xKG Pipeline. (1) Construction: A paper-aware pipeline first curates a corpus of relevant papers and repositories, then extracts, modularizes, and filters it into executable technique-code pairs. (2) Integration: Our xKG guides agents by providing high-level technique overviews and low-level code references.
@@ 2.1 Design Formulation
We model xKG as a hierarchical, multi-relational graph $\mathrm{\textsc{xKG}}=(\mathcal{N},\mathcal{E})$, specifically:
@@
We define three types of nodes to capture knowledge at different granularities:
@@
  • Technique Node ($n_{t}$): A self-contained academic concept $n_{t}=(D_{t},\{n^{\prime}_{t}\}_{k})$ with definition $D_{t}$ and optional sub-nodes $\{n^{\prime}_{t}\}_{k}$, ranging from complete framework to reusable component.
@@
These nodes are linked by the following edges:
@@
  • Implementation Edge ($e_{\text{impl}}$): An edge $(n_{t},n_{c})$ linking a technique to its code implementation.
Our xKG directly links scientific concepts to executable code, yielding more comprehensive knowledge (example in Figure 1(b), Appendix D).
2.2 Paper-Aware xKG Construction
2.2.1 Dynamic Corpus Curation
Our corpus curation is a paper-aware automated pipeline engineered for continuous scalability and cross-domain extensibility, as shown in Figure 1(a). For each paper targeted for reproduction (papers in PaperBench Starace et al. (2025)), xKG automates the collection of all prerequisite resources, beginning by employing o4-mini to identify its core techniques. Centered on these techniques, we first cull its top-ranked references via LLM, and then execute an automated technique-based web retrieval, culminating in a curated set of ten relevant papers for each. Note that we strictly do NOT use any GitHub or third-party reproduction repositories listed in PaperBench’s blacklist to avoid any risk of data leakage. All retrieved papers are processed to fetch their sources from arXiv and then identify the associated GitHub. Papers without official repositories are automatically filtered out, resulting in a corpus of paper-repository pairs.
@@ 2.2.1 Dynamic Corpus Curation
 vanilla +xKG  vanilla +xKG  vanilla +xKG  vanilla +xKG  vanilla +xKG  vanilla +xKG
BasicAgent o3-mini 12.96 $37.22_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+24.26}}$ 22.63 $27.26_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+4.63}}$ 18.24 $20.82_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+2.58}}$ 20.82 $22.86_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+2.04}}$ 14.82 $14.67_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-0.15}}$ 17.89 24.57 $\uparrow$ 6.68
DS-R1 33.05 $39.14_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+6.09}}$ 40.55 $39.14_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-1.41}}$ 17.22 $24.49_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+7.27}}$ 31.56 $33.97_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+2.41}}$ 17.08 $21.38_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+4.30}}$ 27.89 31.62 $\uparrow$ 3.73
IterativeAgent o3-mini 22.22 $43.70_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+21.48}}$ 21.38 $36.28_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+14.90}}$ 28.77 $23.91_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-4.86}}$ 31.28 $29.15_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-2.13}}$ 19.35 $26.50_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+7.15}}$ 24.60 31.91 $\uparrow$ 7.31
DS-R1 16.20 $47.40_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+31.20}}$ 31.19 $31.78_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+0.59}}$ 31.09 $26.57_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-4.52}}$ 35.30 $38.44_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+3.14}}$ 21.32 $31.89_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+10.57}}$ 27.02 35.22 $\uparrow$ 8.20
PaperCoder o3-mini 23.15 $46.48_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+23.33}}$ 45.70 $53.99_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+8.29}}$ 52.48 $52.08_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-0.40}}$ 50.37 $63.13_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+12.76}}$ 39.84 $50.36_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+10.52}}$ 42.31 53.21 $\uparrow$ 10.90
DS-R1 43.24 $49.26_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+6.02}}$ 43.26 $59.19_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+15.93}}$ 51.18 $73.03_{{\color[rgb]{0.15625,0.65625,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.65625,0.26953125}+21.85}}$ 61.12 $60.68_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-0.44}}$ 62.37 $59.53_{{\color[rgb]{0.86328125,0.20703125,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.20703125,0.26953125}-2.84}}$ 52.23 60.34 $\uparrow$ 8.11
@@
Table 1: Main results on PaperBench Code-Dev. We evaluate on the official lite subset of PaperBench (details in Table 4). All results are the best@3 Replication Score (%) to mitigate task stochasticity and potential tool failures.
2.2.2 Hierarchical KG Construction
@@
Based on the corpus, we then construct the xKG with the automated steps shown in Figure 1(b):
@@
  • Step 1: Technique Extraction. We first use o4-mini to deconstruct the paper’s decomposable methodology into a preliminary hierarchical tree of Technique Nodes $\mathcal{N}_{T}$ linked by Structural Edges $e_{\text{struct}}$(details in Appendix B.3). Subsequently, we utilize RAG [^2] We employ text-embedding-3-small for embedding similarity throughout all stages of xKG construction. to enrich each node by retrieving relevant text from the paper to form a comprehensive definition $D_{t}$. This step yields a set of detailed techniques yet may contain noise.
@@
  • Step 2: Code Modularization. For each technique $n_{t}$, its definition is used as a query to retrieve code snippets from the repository via embedding similarity. We then employ o4-mini to synthesize these snippets into candidate Code Nodes $n_{c}$, each includes the implementation $\sigma$, test script $\tau$, and documentation $\delta$. These nodes then undergo an iterative self-debugging loop to ensure executability, ultimately producing a set of executable Code Nodes $\mathcal{N}_{c}$ with associated technique-linking Implementation Edges $e_{\text{impl}}$.
@@
Finally, we construct xKG from 42 curated papers, totaling 591,145 tokens. Our xKG construction is built upon a paper-aware pipeline designed for continuous evolution as new literature emerges, as detailed in Appendix B.1.
2.3 Paper-Guided xKG Integration
xKG can be seamlessly integrated into a practical reproduction workflow, where LLM agents utilize it at two stages (Figure 1, right). For high-level planning, the agent fetches the target paper’s Paper Node (without Code Nodes) to grasp its core techniques and overall structure. During low-level implementation, the agent queries xKG for (Technique, Code) pairs directly relevant to the target paper. These two steps can be supplied either as callable tools for ReAct agents or as pluggable components of fixed-workflow agents. Crucially, all retrieved candidates are processed by a final LLM-based Verifier (o4-mini), acting as a final quality gate to ensure the retrieved knowledge is both technically relevant and implementable.